OpenStudy (anonymous):

Joe Fleet can jog at the rate of 8 mph. Ron Pack can hike carrying camping equipment at the rate of 4 mph. They leave from opposite ends of a 44-mile trail at the same time. How far does each travel if Ron has 15 minutes to set up camp before Joe arrives?

OpenStudy (anonymous):

Joe's travel time in hours (x): 4(x - 0.25) + 8x = 44 4x - 1 + 8x = 44 12x = 45 x = 3.75 Distance covered by Ron: = 4(3.75 - 0.25) = 4(3.5) = 14 Distance covered by Joe: = 8(3.75) = 30 Answer: Ron, 14 miles; Joe, 30 miles

OpenStudy (anonymous):

THANKS VERY MUCH!!!

OpenStudy (anonymous):

np :D